# Learn2Mix: Training Neural Networks Using Adaptive Data Integration

### v2

@@ Abstract
Accelerating model convergence within resource-constrained environments is critical to ensure fast and efficient neural network training. This work presents learn2mix, a novel training strategy that adaptively adjusts class proportions within batches, focusing on classes with higher error rates. Unlike classical training methods that use static class proportions, learn2mix continually adapts class proportions during training, leading to faster convergence. Empirical evaluations conducted on benchmark datasets show that neural networks trained with learn2mix converge faster than those trained with existing approaches, achieving improved results for classification, regression, and reconstruction tasks under limited training resources and with imbalanced classes. Our empirical findings are supported by theoretical analysis.
footnotetext: GitHub repository: https://tinyurl.com/yy7tfsp5.
@@ 1 Introduction
Deep neural networks have become essential tools across various applications of machine learning, including computer vision [Krizhevsky et al., 2012, Simonyan and Zisserman, 2014, He et al., 2016], natural language processing [Vaswani et al., 2017, Devlin et al., 2018, Radford et al., 2019, Touvron et al., 2023], and speech recognition [Hinton et al., 2012, Baevski et al., 2020]. Despite their ability to learn and model complex, nonlinear relationships, deep neural networks often require substantial computational resources during training. In resource-constrained environments, this demand poses a significant challenge [Goyal et al., 2017], making the development of efficient and scalable training methodologies increasingly crucial to fully leverage the capabilities of deep neural networks.
@@
Training deep neural networks relies on the notion of empirical risk minimization [Vapnik and Bottou, 1993], and typically involves optimizing a loss function using gradient-based algorithms [Rumelhart et al., 1986, Bottou, 2010, Kingma and Ba, 2014]. Techniques such as regularization [Srivastava et al., 2014, Ioffe and Szegedy, 2015] and data augmentation [Shorten and Khoshgoftaar, 2019], learning rate scheduling, [Smith, 2017] and early stopping [Prechelt, 1998], are commonly employed to enhance generalization and prevent overfitting. However, the efficiency of the training process itself remains a critical concern, particularly in terms of convergence speed and computational resources.
@@
Within this context, adaptive training strategies, which target enhanced generalization by modifying aspects of the training process, have emerged as promising approaches. Methods such as curriculum learning [Bengio et al., 2009, Graves et al., 2017, Wang et al., 2021] adjust the order and difficulty of training samples to facilitate more effective learning. Insights from these adaptive training strategies can be extended to the class imbalance problem [Wang et al., 2019], where underrepresented classes are intrinsically harder to learn due to data scarcity [Buda et al., 2018], a challenge intensified in adversarial settings where safe data collection is severely limited [Wang and Gursoy, 2023]. These methods are typically categorized into data-level methods, such as oversampling and undersampling [Chawla et al., 2002], and algorithm-level approaches, including class-balanced loss functions [Lin et al., 2017]. However, developing adaptive neural network training methodologies that accelerate model convergence, while ensuring robustness to class imbalance, remains an open problem.
@@
Building upon these insights, a critical aspect of training efficiency lies in the composition of batches used during stochastic gradient descent. Classical training paradigms maintain approximately fixed class proportions within each shuffled batch, mirroring the overall class distribution in the training dataset [Buda et al., 2018, Peng et al., 2019]. However, this static approach fails to account for the varying levels of difficulty associated with different classes, which can hinder optimal convergence rates. For example, classes with higher error rates or those that are inherently more challenging may require greater emphasis during training to enhance model performance. Ignoring these nuances can lead to suboptimal learning trajectories and prolonged training periods. While existing approaches address class imbalance by adjusting sample weights or dataset resampling, they do not dynamically change the class-wise composition of batches during training via real-time performance metrics.
@@
This observation motivates the central question of this paper: Can we dynamically adjust the proportion of classes within batches, across training epochs, to accelerate model convergence? Addressing this question involves developing strategies that dynamically modify the proportion of classes using real-time performance metrics, thereby directing the learning procedure towards more challenging or underperforming classes. Such adaptive batch construction has the potential to enhance convergence rates and model accuracy, providing more efficient training, especially within scenarios characterized by class imbalance or heterogeneous class difficulties [Liu et al., 2008, Ren et al., 2018].
To address these considerations, in this work, we introduce learn2mix, a novel training strategy that dynamically modifies class proportions in batches by emphasizing classes with higher instantaneous error rates. In contrast with classical training schemes that utilize fixed class proportions, learn2mix continually adapts these proportions during training via real-time class-wise error metrics. This dynamic adjustment facilitates faster convergence and improved performance across various tasks, including classification, regression, and reconstruction. An illustration of the learn2mix training methodology is provided in Figure 1, demonstrating the adaptive class-wise composition of batches.
@@
    We propose learn2mix, a novel adaptive training strategy that dynamically adjusts class proportions within batches, utilizing class-wise error rates, to accelerate model convergence in resource constrained regimes.
@@
    We prove that neural networks trained using learn2mix converge faster than those trained using classical approaches when certain properties hold, wherein the class proportions converge to a stable distribution proportional to the optimal class-wise error rates.
@@
The landscape of neural network training methods is characterized by a diverse set of approaches aiming to enhance model performance and training efficiency. Handling class imbalance has been extensively analyzed, with methods including oversampling [Chawla et al., 2002], undersampling [Tahir et al., 2012], and class-balanced loss functions [Lin et al., 2017, Ren et al., 2018] being proposed to mitigate biases towards majority classes. In parallel, curriculum learning [Bengio et al., 2009] and reinforcement learning-centric approaches [Florensa et al., 2017] have introduced methods to facilitate more effective learning trajectories. Meta-learning, or learn2learn methodologies [Arnold et al., 2020], including model-agnostic meta-learning (MAML) [Finn et al., 2017], focus on optimizing the learning process itself to enable rapid adaptation to new tasks, highlighting the importance of adaptability in model training. Additionally, adaptive data sampling strategies [Liu et al., 2008] and boosting algorithms [Freund and Schapire, 1997] emphasize the significance of prioritizing harder or misclassified examples to improve model robustness and convergence rates. Despite these advances, most existing training approaches either adjust sample weights, resample datasets, or modify the sequence of training examples without specifically altering the class proportions within batches in an adaptive manner. Our proposed learn2mix strategy distinguishes itself by continually adapting class proportions in these batches throughout the training process, directly targeting classes with higher error rates to accelerate convergence. This approach characterizes a unified framework by addressing class imbalance and integrating adaptive training principles, enhancing training efficiency.
@@ 2 Theoretical Results
[…]
@@
Consider $\tilde{\alpha}\in[0,1]^{k}$ as the vector of fixed-proportion mixing parameters, and let $\mathcal{L}(\theta^{t})\in\mathbb{R}^{k}$ be the vector of class-wise losses at time $t$. The loss for classical training at time $t$ is:
@@
Consider $\alpha^{t},\alpha^{t-1}\in[0,1]^{k}$ as the vector of mixing parameters at time $t$ and time $t-1$, and let $\mathcal{L}(\theta^{t}),\mathcal{L}(\theta^{t-1})\in\mathbb{R}^{k}$ denote the respective class-wise loss vectors at time $t$ and time $t-1$. Consider $\gamma\in(0,1)$ as the mixing rate. The loss for learn2mix training at time $t$ is:
@@
[…]
@@
$\theta^{t+1}=\theta^{t}-\eta\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t}),\quad\quad\text{with: \quad$\eta\in\mathbb{R}_{>0}$},$ (4)
@@
$\lim_{t\rightarrow\infty}\theta^{t}=\theta^{*},\quad\ \ \text{and:}\ \ \lim_{t\rightarrow\infty}\alpha^{t}=\alpha^{*}=\frac{\mathcal{L}(\theta^{*})}{\mathds{1}_{k}^{T}\mathcal{L}(\theta^{*})}.$ (5)
@@
Let $\mathcal{L}(\theta^{t}),\mathcal{L}(\theta^{*})\in\mathbb{R}^{k}$ denote the respective class-wise loss vectors for the model parameters at time $t$ and for the optimal model parameters. Suppose each class-wise loss, $\mathcal{L}_{i}(\theta)\in\mathbb{R}$ is strongly convex in $\theta$ with strong convexity parameter $\mu_{i}\in\mathbb{R}_{>0}$, $\forall i\in\{1,\ldots,k\}$, and each class-wise loss gradient $\smash{\nabla_{\theta}\mathcal{L}_{i}(\theta)\in\mathbb{R}^{m}}$ is Lipschitz continuous in $\theta$, having Lipschitz constant $L_{i}\in\mathbb{R}_{\geq 0}$, $\forall i\in\{1,\ldots,k\}$. Moreover, suppose that the loss gradient $\smash{\nabla_{\theta}\mathcal{L}(\theta,\alpha)\in\mathbb{R}^{m}}$ is Lipschitz continuous in $\alpha$, having Lipschitz constant $L_{\alpha}\in\mathbb{R}_{\geq 0}$, and let $\smash{\mu^{*}=\min_{i\in\{1,\ldots,k\}}\mu_{i}}$, $\smash{L^{*}=\max_{i\in\{1,\ldots,k\}}L_{i}}$. Then, iff the following holds:
$\begin{split}&\Big{[}\Big{(}\frac{\mu^{*}}{2}-L^{*}\Big{)}\|\theta^{t}-\theta^{*}\|^{2}+\tilde{\alpha}^{T}(\mathcal{L}(\theta^{t})-\mathcal{L}(\theta^{*}))\Big{]}\\ &\quad\quad\ \times\Big{[}\|\theta^{t}-\theta^{*}\|-(\mathcal{L}(\theta^{t})-\mathcal{L}(\theta^{*}))\Big{]}>0,\end{split}$ (8)
@@
$\left\|\left(\theta^{t}\,\text{--}\,\eta\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})\right)\text{--}\ \theta^{*}\right\|\leq\left\|\left(\theta^{t}\,\text{--}\,\eta\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha})\right)\text{--}\ \theta^{*}\right\|$ (9)
The formula for $\beta$ can be found in Section A of the Appendix.
@@ 3 Algorithm
$\qquad\qquad\qquad\mathcal{L}(\theta^{t},\alpha^{t})=\sum_{i=1}^{k}\alpha_{i}^{t}\mathcal{L}_{i}(\theta^{t})$ (11)
$\quad=\sum_{i=1}^{k}\alpha_{i}^{t}\bigg{[}\frac{1}{P}\sum_{p=1}^{P}\bigg{[}\frac{1}{\alpha_{i}^{t}M}\sum_{(x_{j},y_{j})\in S_{i}^{p}}\ell(h_{\theta^{t}}(x_{j}),y_{j})\bigg{]}\bigg{]}.$
@@
3       $\alpha_{i}\leftarrow\tilde{\alpha}_{i}$   (Initialize time-varying mixing params)

### v3

@@ Abstract
Accelerating model convergence in resource-constrained environments is essential for fast and efficient neural network training. This work presents learn2mix, a new training strategy that adaptively adjusts class proportions within batches, focusing on classes with higher error rates. Unlike classical training methods that use static class proportions, learn2mix continually adapts class proportions during training, leading to faster convergence. Empirical evaluations on benchmark datasets show that neural networks trained with learn2mix converge faster than those trained with existing approaches, achieving improved results for classification, regression, and reconstruction tasks under limited training resources and with imbalanced classes. Our empirical findings are supported by theoretical analysis.
footnotetext: GitHub repository: https://github.com/shyamven/Learn2Mix.
@@ 1 Introduction
Deep neural networks have become essential tools across various applications of machine learning, including computer vision (Krizhevsky et al., 2012; Simonyan and Zisserman, 2014; He et al., 2016), natural language processing (Vaswani et al., 2017; Devlin et al., 2018; Radford et al., 2019; Touvron et al., 2023), and speech recognition (Hinton et al., 2012; Baevski et al., 2020). Despite their ability to learn and model complex, nonlinear relationships, deep neural networks often require substantial computational resources during training. In resource-constrained environments, this demand poses a significant challenge (Goyal et al., 2017), making the development of efficient and scalable training methodologies increasingly crucial to fully leverage the capabilities of these models.
@@
Training deep neural networks relies on the notion of empirical risk minimization (Vapnik and Bottou, 1993), and typically involves optimizing a loss function using gradient-based algorithms (Rumelhart et al., 1986; Bottou, 2010; Kingma and Ba, 2014). Techniques such as regularization (Srivastava et al., 2014; Ioffe and Szegedy, 2015) and data augmentation (Shorten and Khoshgoftaar, 2019), learning rate scheduling, (Smith, 2017) and early stopping (Prechelt, 1998), are commonly employed to enhance generalization and prevent overfitting. However, the efficiency of the training process itself remains a critical concern, particularly in terms of convergence speed and computational resources.
@@
Within this context, adaptive training strategies, which target enhanced generalization by modifying aspects of the training process, have emerged as promising approaches. Methods such as curriculum learning (Bengio et al., 2009; Graves et al., 2017; Wang et al., 2021) adjust the order and difficulty of training samples to facilitate more effective learning. Insights from these adaptive training strategies can be extended to the class imbalance problem (Wang et al., 2019), where underrepresented classes are intrinsically harder to learn due to data scarcity (Buda et al., 2018), a challenge intensified in adversarial settings where safe data collection is severely limited (Wang and Gursoy, 2023). These methods are typically categorized into data-level methods, such as oversampling and undersampling (Chawla et al., 2002) and algorithm-level schemes, including class-balanced loss functions (Lin et al., 2017). However, developing adaptive neural network training methodologies that accelerate model convergence, while ensuring robustness to class imbalance, remains an open problem.
@@
Building upon these insights, a critical aspect of training efficiency lies in the composition of batches used during stochastic gradient descent. Classical training paradigms maintain approximately fixed class proportions within each shuffled batch, mirroring the overall class distribution in the training dataset (Buda et al., 2018; Peng et al., 2019). However, this static approach fails to account for the varying levels of difficulty associated with different classes, which can hinder optimal convergence rates. For example, classes with higher error rates or those that are inherently more challenging may require greater emphasis during training to enhance model performance. While existing approaches address class imbalance by adjusting sample weights or dataset resampling, they do not dynamically change the class-wise composition of batches during training via real-time performance metrics.
@@
This observation motivates the central question of this paper: Can we adaptively adjust the proportion of classes within batches, across training epochs, to accelerate model convergence? Addressing this question involves developing approaches that dynamically modify class proportions using real-time performance metrics, directing learning towards underperforming classes. Such batch construction has the potential to enhance convergence rates and training efficiency, especially in scenarios with imbalanced classes or heterogeneous class difficulties (Liu et al., 2008; Ren et al., 2018).
To address these considerations, in this work, we introduce learn2mix, a novel training strategy that dynamically modifies class proportions in batches by emphasizing classes with higher instantaneous error rates. In contrast with classical training schemes that utilize fixed class proportions, learn2mix continually adapts these proportions during training via real-time class-wise error rates. This dynamic adjustment facilitates faster convergence and improved performance across various tasks, including classification, regression, and reconstruction. An illustration of the learn2mix training methodology is provided in Figure 1, demonstrating the adaptive class-wise composition of batches.
@@
    We propose learn2mix, a novel adaptive training strategy that dynamically adjusts class proportions within batches, utilizing class-wise error rates, to accelerate model convergence.
@@
    We prove that neural networks trained using learn2mix converge faster than those trained using classical approaches when certain properties hold, such that the class proportions converge to a stable distribution proportional to the optimal class-wise error rates.
@@
The landscape of neural network training methods comprises various approaches aiming to enhance model performance and training efficiency. Handling class imbalance has been extensively studied, with methods such as importance sampling (Katharopoulos and Fleuret, 2018), oversampling (Chawla et al., 2002), undersampling (Tahir et al., 2012), and class-balanced loss functions (Lin et al., 2017; Ren et al., 2018) being proposed to mitigate biases towards majority classes. In parallel, curriculum learning (Bengio et al., 2009) and reinforcement learning approaches (Florensa et al., 2017) have introduced methods to facilitate more effective learning trajectories. Meta-learning, or learn2learn methodologies (Arnold et al., 2020), including model-agnostic meta-learning (MAML) (Finn et al., 2017), focus on optimizing the learning process itself to enable rapid adaptation to new tasks. Additionally, adaptive data sampling strategies (Liu et al., 2008) and boosting algorithms (Freund and Schapire, 1997) emphasize the significance of prioritizing harder or misclassified examples to improve model robustness. Despite these advances, most existing training methods either adjust sample weights, resample datasets, or modify the sequence of training examples without specifically altering the class proportions within batches in an adaptive manner. Our proposed learn2mix strategy distinguishes itself by adapting batch class proportions throughout the training process, targeting classes with higher error rates to accelerate convergence. This approach offers a unified framework by addressing class imbalance through adaptive training principles.
@@ 2 Theoretical Results
Consider the random variables $X\in\mathbb{R}^{d}$ and $Y\in\mathbb{R}^{k}$, where $X$ denotes the feature vector, $Y$ are the labels, and $k$ is the number of classes. We consider the original training dataset, $J=\{(x_{j},y_{j})\}_{j=1}^{N}$, where $\smash{(x_{j},y_{j})\overset{\text{i.i.d.}}{\sim}(X,Y)}$, $\forall j\in\{1,\ldots,N\}$. The class proportions for this dataset are given by the vector of fixed-proportion mixing parameters, $\tilde{\alpha}=[\tilde{\alpha}_{1},\ldots,\tilde{\alpha}_{k}]^{T}$, reflecting the distribution of classes. We define $\alpha=[\alpha_{1},\ldots,\alpha_{k}]^{T}$ as a variable denoting the vector of mixing parameters, where $\alpha_{i}\in[0,1]$ and $\smash{\sum_{i=1}^{\raisebox{-0.75346pt}{$\scriptstyle k$}}\alpha_{i}=1}$. The value of $\alpha$ determines the class proportions used during training, and can vary depending on the chosen training mechanism. In classical training, $\alpha=\alpha^{t}$ is constant over time and reflects the class proportions within the original training dataset, wherein $\smash{\alpha^{t}=\tilde{\alpha}}$, $\forall t\in\mathbb{N}$. In learn2mix training, $\alpha=\alpha^{t}$ is time-varying, and is initialized at time $t=0$ as $\alpha^{0}=\tilde{\alpha}$.
@@
Consider $\tilde{\alpha}\in[0,1]^{k}$ as the vector of fixed-proportion mixing parameters, and let $\mathcal{L}(\theta^{t})\in\mathbb{R}^{k}$ denote the vector of class-wise losses at time $t$. The loss for classical training at time $t$ is given by:
@@
Consider $\alpha^{t},\alpha^{t-1}\in[0,1]^{k}$ as the vector of mixing parameters at time $t$ and time t1t-1, and let $\mathcal{L}(\theta^{t}),\mathcal{L}(\theta^{t-1})\in\mathbb{R}^{k}$ denote the respective class-wise loss vectors at time $t$ and time t1t-1. Consider $\gamma\in(0,1)$ as the mixing rate. The loss for learn2mix training at time $t$ is given by the following:
@@
$\text{Where: }\alpha^{t}=\alpha^{t-1}+\gamma\bigg(\frac{\mathcal{L}(\theta^{t-1})}{\mathds{1}_{k}^{T}\mathcal{L}(\theta^{t-1})}-\alpha^{t-1}\bigg),$ (3)
We note that the denominator, $\mathds{1}_{k}^{T}\mathcal{L}(\theta^{t-1})$, is the sum of losses across all classes, and dividing by it converts $\mathcal{L}(\theta^{t-1})$ into a probability distribution. We update $\alpha^{t-1}$ by nudging the mixing parameters toward this probability distribution, so classes with higher losses receive a larger share of samples in the next time step. The scalar mixing rate, $\gamma$, is a user-defined step size hyperparameter that controls how aggressively $\alpha^{t-1}$ moves. We note that classical training is recovered by setting $\gamma=0$.
@@
Let $\mathcal{L}(\theta^{t}),\mathcal{L}(\theta^{*})\in\mathbb{R}^{k}$ denote the respective class-wise loss vectors for the model parameters at time $t$ and for the optimal model parameters. Suppose each class-wise loss $\mathcal{L}_{i}(\theta)\in\mathbb{R}$ is strongly convex in $\theta$, with strong convexity parameter $\mu_{i}\in\mathbb{R}_{>0}$, $\forall i\in\{1,\ldots,k\}$, and each class-wise loss gradient $\smash{\nabla_{\theta}\mathcal{L}_{i}(\theta)\in\mathbb{R}^{m}}$ is Lipschitz continuous in $\theta$, having Lipschitz constant $L_{i}\in\mathbb{R}_{\geq 0}$, $\forall i\in\{1,\ldots,k\}$. Let $\smash{\mu^{*}=\min_{i\in\{1,\ldots,k\}}\mu_{i}}$, $\smash{L^{*}=\max_{i\in\{1,\ldots,k\}}L_{i}}$. Then, if the model parameters at time $t+1$ are obtained via the gradient of the loss for learn2mix training, where:
$\theta^{t+1}=\theta^{t}-\eta\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t}),\quad\quad\text{with: \qquad$\eta\in\mathbb{R}_{>0}$},$ (4)
@@
Let $\mathcal{L}(\theta^{t})\in\mathbb{R}^{k}$ denote the class-wise loss vector at time $t$. Suppose each class-wise loss, $\mathcal{L}_{i}(\theta)\in\mathbb{R}$, is strongly convex in $\theta$, with strong convexity parameter $\mu_{i}\in\mathbb{R}_{>0}$, $\forall i\in\{1,\ldots,k\}$, and suppose each class-wise loss gradient $\smash{\nabla_{\theta}\mathcal{L}_{i}(\theta)\in\mathbb{R}^{m}}$ is Lipschitz continuous in $\theta$ with Lipschitz constant $L_{i}\in\mathbb{R}_{\geq 0}$, $\forall i\in\{1,\ldots,k\}$. Let $\smash{\mu^{*}=\min_{i\in\{1,\ldots,k\}}\mu_{i}}$, $\smash{L^{*}=\max_{i\in\{1,\ldots,k\}}L_{i}}$. Then, the following condition and inequality hold, $\forall\alpha\in[0,1]^{k}$ where $\smash{\sum_{i=1}^{\raisebox{-0.75346pt}{$\scriptstyle k$}}\alpha_{i}=1}$:
@@
$\text{Wherein: }\|\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})\|+\|\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha})\|\leq 2L^{*}\|\theta^{t}-\theta^{*}\|.$ (7)
@@
Let $\mathcal{L}(\theta^{t}),\mathcal{L}(\theta^{*})\in\mathbb{R}^{k}$ denote the respective class-wise loss vectors for the model parameters at time $t$ and for the optimal model parameters. Suppose each class-wise loss, $\mathcal{L}_{i}(\theta)\in\mathbb{R}$ is strongly convex in $\theta$ with strong convexity parameter $\mu_{i}\in\mathbb{R}_{>0}$, $\forall i\in\{1,\ldots,k\}$, and each class-wise loss gradient $\smash{\nabla_{\theta}\mathcal{L}_{i}(\theta)\in\mathbb{R}^{m}}$ is Lipschitz continuous in $\theta$, having Lipschitz constant $L_{i}\in\mathbb{R}_{\geq 0}$, $\forall i\in\{1,\ldots,k\}$. Moreover, suppose the loss gradient $\smash{\nabla_{\theta}\mathcal{L}(\theta,\alpha)\in\mathbb{R}^{m}}$ is Lipschitz continuous in $\alpha$, having Lipschitz constant $L_{\alpha}\in\mathbb{R}_{\geq 0}$, and let $\smash{\mu^{*}=\min_{i\in\{1,\ldots,k\}}\mu_{i}}$, $\smash{L^{*}=\max_{i\in\{1,\ldots,k\}}L_{i}}$. Then, if and only if the following condition holds:
@@
The complete formula for $\beta$ can be found in Section A of the Appendix.
@@ 3 Algorithm
$\mathcal{L}(\theta^{t},\alpha^{t})=\sum_{i=1}^{k}\alpha_{i}^{t}\mathcal{L}_{i}(\theta^{t})=\sum_{i=1}^{k}\alpha_{i}^{t}\bigg[\frac{1}{P}\sum_{p=1}^{P}\bigg[\frac{1}{\alpha_{i}^{t}M}\sum_{(x_{j},y_{j})\in S_{i}^{p}}\ell(h_{\theta^{t}}(x_{j}),y_{j})\bigg]\bigg].$ (11)
@@ 2 $J_{i}\leftarrow\{(x_{j},y_{j})\}_{j=1}^{\alpha_{i}N}$ (Initialize class-specific training datasets)
3 $\alpha_{i}\leftarrow\tilde{\alpha}_{i}$ (Initialize time-varying mixing parameters)